A capillary tube is immersed in water of surface tension 70 x 10-2 Nm-1 and rises 6.2 cm. By what depth will mercury be depressed if the same capillary is immersed in it? (Surface tension of mercury = 0.54 Nm-1; angle of contact between mercury and glass = 140 degrees, density of mercury = 13600 kgm-3)
1
Expert's answer
2020-10-23T11:54:32-0400

As the height of liquid rise in a capillary is equal to: "h=\frac{2\times σ\times \cos \theta}{p\times g\times R}" , where: "\u03c3" -surface tension coefficient, "p" is the fluid density, "g" -acceleration of free fall, "R" is the radius of the capillary.

Then: "R= \frac{p\times g\times h} {2\times σ\times \cos \theta }= \frac{10^3\times 10\times 0.062} {2\times 0.7\times \cos 0 }\approx442.86" m.

Hence the height at mercury: "h=\frac{2\times σ\times \cos \theta}{p\times g\times R}=" "\frac{2\times 0.54\times \cos {(180\degree-140\degree)}}{13600\times 10\times 442.86}\approx1. 37\times 10^{-8}" m.